OpenStudy (anonymous):

How to simplify: 6√v - √4v^2 - √36v + √v^2 ? I can't find the right answer, but this is what I have written down.. I don't believe the first one can be simplified -√4v^2 = -√2 * v^2 = -√2 * √v^2 = -2√v^2 -√36v = -√6 * v = -√6 * √v = -6√v I don't think the last one can be simplified. So when I combine everything I have: 6√v - 2√v^2 - 6√v + √v^2 6√v - 6√v = 0 - 2√v^2 + √v^2 = -1√v^2 But the only answers that I have to choose from are: 3v v√v -v 6v^2 I'm not sure what I've done wrong.

OpenStudy (phi):

first is the 2nd term \[ - \sqrt{4v^2} \] with the v^2 under the radical ?

OpenStudy (phi):

but in general you should remember that \[ \sqrt{x^2} = x \] or \[ \sqrt{x\cdot x}= x \]

OpenStudy (anonymous):

\[6\sqrt{v}- \sqrt{4v^{2}}-\sqrt{36v}+\sqrt{v^{2}}\] \[6\sqrt{v}- 2v-6\sqrt{v}+v\]

OpenStudy (phi):

so right away you should be able to simplify the last term

OpenStudy (anonymous):

But since the last term is √v^2, wouldn't you just add that with -2√v^2?

OpenStudy (anonymous):

yes and thats how u ll arrive at the final answer

OpenStudy (anonymous):

Oh okay, so then it would be -v?

OpenStudy (anonymous):

Yes!